Show that: (a) det(A) - (b) det (A2) [det (A)2
Answer #1

(a)

We know that

det(A . B) = det(A) . det(B)

and A. A-1 = I

where I is the unit matrix.

For a unit matrix I, det(I) = 1

So, det(A. A-1 ) = det(I)

=> det(A) . det(A-1) = 1

=> det(A) = 1 / det(A-1)

(b)

We know that det(A . B) = det(A) . det(B)

Let A = B, then det(A . A) = det(A) . det(A)

Now,

det(A2) = det(A . A) = det(A) . det(A) = [det(A)]2

Thus, det(A2) = [det(A)]2